OpenStudy (anonymous):

If cos(x) = A, then cos (x + π)= A) A B) -A C) 2A Eliminate D) A + π

OpenStudy (anonymous):

itsD right?

OpenStudy (anonymous):

@Fandude727

OpenStudy (anonymous):

cos(x + pi) = a + pi?

OpenStudy (fandude727):

does this help? http://www.wolframalpha.com/input/?i=cos(x)+%3D+sin(x%2Bpi%2F2)

OpenStudy (anonymous):

Its -A since pi = 180 i figured it out

OpenStudy (tkhunny):

Let's see. \(\cos(A+\pi) = \cos(A)\cos(\pi) - \sin(\pi)\sin(A) = \cos(A)(-1) - (0)\sin(A)\) Are we getting anywhere?

OpenStudy (anonymous):

Dude i already figured it it, -a was correct since pi = a 180 degree change making a go into the next quadrant causing it to be in the negative quadrant